# arXiv:2603.02804v1[quant-ph] 03 Mar 2026

### 03 Mar 2026 v1

@@ Abstract
While real quantum devices have been increasingly used to conduct research focused on achieving quantum advantage or quantum utility in recent years, executing deep quantum circuits or performing quantum machine learning with large-scale data on current noisy intermediate-scale quantum devices remains challenging, making classical simulation essential for quantum machine learning research. However, classical simulation often suffers from the cost of gradient calculations, requiring enormous memory or computational time. In this paper, to address these problems, we propose a method to fuse multiple consecutive gates in each of the forward and backward paths to improve throughput by minimizing global memory accesses. As a result, we achieved approximately $20$ times throughput improvement for a Hardware-Efficient Ansatz with $12$ or more qubits, reaching over $30$ times improvement on a mid-range consumer GPU with limited memory bandwidth. By combining our proposed method with gradient checkpointing, we drastically reduce memory usage, making it possible to train a large-scale quantum machine learning model, a $20$-qubit, $1,000$-layer model with $60,000$ parameters, using $1,000$ samples in approximately $20$ minutes. This implies that we can train the model on large datasets, consisting of tens of thousands of samples, such as MNIST or CIFAR-10, within a realistic time frame (e.g., $20$ hours per epoch). In this way, our proposed method drastically accelerates classical simulation of quantum machine learning, making a significant contribution to quantum machine learning research and variational quantum algorithms, such as verifying algorithms on large datasets or investigating learning theories of deep quantum circuits like barren plateau.
@@ 1 Introduction
A popular approach for gradient calculation is the parameter-shift rule [mitarai2018quantum]. This approach is memory-efficient, but the number of quantum circuit executions scales linearly with the number of variational gates, making it impractical for large-scale models in classical simulation. On the other hand, the adjoint method is fast because the effective number of quantum circuit executions is independent of the number of variational gates. However, a practical implementation of the adjoint method faces a critical trade-off between memory usage and execution time regarding the management of intermediate states. A naive implementation stores all intermediate states of quantum gates in the forward path in order to reuse them in the backward path, causing memory usage to increase linearly with the number of gates. Alternatively, a memory-efficient implementation avoids storing intermediate states by sequentially applying each gate to recompute them during the backward path [jones2020efficient, luo2020yao], leading to a significant increase in computational time due to the repeated accesses to exponentially large quantum states [jones2020efficient, luo2020yao]. Therefore, existing simulators typically optimize only forward paths using gate fusion [suzuki2021qulacs, haner20175, smelyanskiy2016qhipster], or sequentially apply quantum gates to obtain intermediate states in the backward paths [jones2020efficient, luo2020yao], leaving the backward path bottlenecked by inefficient memory accesses.
@@
In this paper, we propose a method to fuse multiple consecutive quantum gates into a single fused operator for each of the forward and backward paths, without writing any intermediate states of the fused gates to global memory. This approach allows us to reduce the memory requirements for storing quantum states and decrease the number of accesses to global memory, leading to an enhancement of the arithmetic intensity. However, if we simply read the input state and temporarily store all recomputed intermediate states, the number of variables to be held in registers can easily exceed the register capacity. This causes register spilling and increases the computation time. Therefore, to reduce this register pressure, we minimize the number of intermediate states held in registers, and load unitary matrices whenever necessary. As a result, our proposed method achieves approximately $20$ times improvement in throughput compared to a PyTorch native implementation for a typical single-layer Hardware-Efficient Ansatz (HEA) [kandala2017hardware]. In particular, for a mid-range GPU with limited memory bandwidth, we achieved over $30$ times improvement in throughput, highlighting the effectiveness of our proposed method on GPUs with limited memory bandwidth.
@@
Moreover, since our implementation uses Triton [triton2019] and PyTorch [Ansel_PyTorch_2_Faster_2024] and is integrated with PyTorch’s automatic differentiation, it is easy to utilize the existing PyTorch ecosystem. For example, by combining our proposed method with gradient checkpointing, we can reduce memory usage to $O(\sqrt{d})$ for $d$ HEA layers. In our numerical experiment, we performed a task of minimizing a sum of batched expectation values for an observable at the final states of a $20$-qubit with $1,000$ HEA layers ($60,000$ parameters). As a result, we succeeded in running $1,000$ samples in approximately $20$ minutes per epoch. This result indicates that we can train a model consisting of $60,000$ parameters using tens of thousands of samples within $20$ hours per epoch, i.e., a realistic time frame. Further acceleration of training can be easily realized using multiple GPUs with PyTorch’s distributed training utilities, such as DistributedDataParallel.
@@ 2.2 Adjoint method
Note that from Eq. (3), we need both the input quantum state $|\psi_{j-1}\rangle$ and the adjoint state $|\lambda_{j}\rangle$ to calculate the gradient with respect to each trainable parameter $\theta_{j}$. This indicates that we need to preserve the input quantum states $|\psi_{j-1}\rangle$ to efficiently compute the gradients. Namely, a naive implementation requires $O(M\cdot 2^{n})$ memory to store all the input quantum states $\{|\psi_{0}\rangle,\dots,|\psi_{M-1}\rangle\}$ in memory. As the number of variational parameters $M$ grows, the required memory can easily exceed the memory capacity, making classical simulation infeasible. Even within the memory capacity, the computation time significantly increases due to repeated accesses to these exponentially large vectors in global memory. We will tackle these problems with our proposed method in the following sections.
@@ 3.1 Applying a quantum gate
As we mentioned in Sec. 2.1, an $n$-qubit state vector $|\psi\rangle$ is represented by a complex vector of length $2^{n}$. We describe $|\psi\rangle$ using computational basis states $|k\rangle$ and complex probability amplitudes $a_{k}$, whose index $k$ $(0\leq k<2^{n})$ corresponds to its binary representation $b_{n-1}\ldots b_{0}$ $(k=\sum_{i=0}^{n-1}b_{i}2^{i})$ of length $n$, as follows:
$|\psi\rangle=\sum_{k=0}^{2^{n}-1}a_{k}|k\rangle=a_{0\ldots 00}|0\cdots 00\rangle+a_{0\ldots 01}|0\cdots 01\rangle+\cdots+a_{1\ldots 10}|1\cdots 10\rangle+a_{1\ldots 11}|1\cdots 11\rangle$
@@
Next, we explain how to update amplitudes when applying two-qubit gates. First, when applying a CZ gate acting on control qubit $c$ and target qubit $t$ to a quantum state, we only need to multiply the amplitudes $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$) by 1-1, where the control and target bits of the index are both $1$.
Similarly, when applying CNOT gate acting on control qubit $c$ and target qubit $t$ to a quantum state, we can update the quantum state by swapping the pair of amplitudes $a_{\ldots 1_{c}\ldots 0_{t}\ldots}$ and $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or $a_{\ldots 0_{t}\ldots 1_{c}\ldots}$ and $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$), i.e., swapping the amplitudes corresponding to the $t$-th bit being 0 and $1$ wherever the $c$-th bit is $1$.
@@ 3.2 Gate fusion in the forward path
Since the adjoint method requires the input state vector for each variational gate, our implementation stores only the output state vector of each fused variational gate during the forward path. As detailed in the next subsection, we then use this stored output state vector to recompute the input state vector for each constituent gate within the fused operator.
@@ 3.3 Gate fusion in the backward path
As shown in Eq. (3), when we calculate the gradients of variational parameters $\theta_{j+k}$, we need both the input quantum state $|\psi_{j+k-1}\rangle$ and the adjoint state $|\lambda_{j+k}\rangle$ for the corresponding variational gate. If we simply fuse $m$ consecutive variational gates in the backward path, it requires storing the elements of quantum states $|\psi_{j}\rangle,\ldots,|\psi_{j+m-1}\rangle$, $|\lambda_{j}\rangle,\ldots,|\lambda_{j+m}\rangle$ in registers. Therefore, as $m$ increases, the required memory can easily exceed the register capacity of the GPU, causing them to be spilled to a low-speed memory region, which increases computation time.
Therefore, by focusing on the reversibility of quantum gates, we recompute the intermediate states when needed, instead of storing all of them. Specifically, as shown in Fig. 1, given the quantum state $|\psi_{j+m}\rangle$ (stored during the forward path) and the adjoint state $|\lambda_{j+m}\rangle$ as inputs for the fused gates consisting of $m$ consecutive variational gates, we perform the following procedure to calculate the gradients.
  • Recomputing the input quantum states for $U_{j+k}(\theta_{j+k})$ and calculating the gradients $\frac{\partial L}{\partial\theta_{j+k}}$ for $k=1,\ldots,m$
    We calculate the gradients, input quantum states, and adjoint states for $U_{j+k}(\theta_{j+k})$ by iterating the gate index of the fused gates in reverse order from $k=m$ to $k=1$. For each step $k$,
@@
      We load $U_{j+k}(\theta_{j+k})$ and apply it to $|\psi_{j+k}\rangle$ in order to calculate $|\psi_{j+k-1}\rangle$:
      $|\psi_{j+k-1}\rangle=U_{j+k}^{\dagger}(\theta_{j+k})|\psi_{j+k}\rangle$
@@
      According to Eq. (3), we calculate the gradient $\frac{\partial L}{\partial\theta_{j+k}}$.
@@
      We update the adjoint state using the loaded gate $U_{j+k}(\theta_{j+k})$:
      $|\lambda_{j+k-1}\rangle=U_{j+k}^{\dagger}(\theta_{j+k})|\lambda_{j+k}\rangle$
@@
Using this procedure, each thread only needs to store a part of the elements of two exponentially large vectors ($|\psi_{j+k-1}\rangle$ and $|\lambda_{j+k}\rangle$) in registers during kernel processing. To avoid register spilling, we read $U_{j+k}(\theta_{j+k})$ when required. Compared to the method of storing all the input quantum states and adjoint states, although our method requires additional calculations to recompute the input quantum states for each $U_{j+k}(\theta_{j+k})$, it allows us to mitigate the bottleneck, i.e., decrease the accesses to global memory. This is expected to reduce the required memory and improve throughput by reducing the number of stored quantum states.
@@ 3.4 Theoretical memory analysis with gradient checkpointing
To theoretically analyze the behavior of memory usage, we consider a typical deep parameterized quantum circuit consisting of a repeating layer structure, such as a Hardware-Efficient Ansatz (HEA). Let $d$ be the total number of HEA layers in a quantum circuit, and let $l_{\text{var}}$ and $l_{\text{const}}$ represent the number of variational gates and constant gates per layer, respectively.
In the naive implementation of the adjoint method without gradient checkpointing, since we need to store the input quantum states of all gates for the backward path, the number of required state vectors to be stored is $(l_{\text{var}}+l_{\text{const}})d$. To mitigate this linear memory growth, we apply gradient checkpointing by dividing the quantum circuit into checkpoint blocks, each consisting of $b$ consecutive HEA layers, as illustrated in Fig. 2. For simplicity, we assume that $d$ is divisible by $b$.
When combining the naive implementation with gradient checkpointing, the number of required state vectors to be stored for the backward path becomes $(l_{\text{var}}+l_{\text{const}})b+d/b$. The first term represents the stored intermediate states within a single active checkpoint block, and the second term ($d/b$) arises from the stored input quantum states for each checkpoint block, which are required to recalculate the intermediate state vectors within the checkpoint block during gradient checkpointing. Therefore, letting $M_{\text{sv}}$ be the required memory for a single state vector, the required memory $M_{total}$ (bytes) for the PyTorch native implementation is estimated as
@@
Similarly, for our proposed method combined with gradient checkpointing, we must store $d/b$ input quantum states for the checkpoint blocks. Additionally, within the active checkpoint block, we only need to store $\left\lceil\frac{l_{\text{var}}}{m}\right\rceil\cdot b$ output state vectors of the fused variational gates, since the implementation of our proposed method does not require storing any input and output state vectors for constant gates. By introducing $\alpha$ as an effective number of fused gates to uniformly treat the memory reduction effects regardless of the type of saved tensors, the required memory (bytes) can be estimated as
@@
where the equality holds when $b^{*}=\sqrt{d/\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil}$. Specifically, for our proposed method (Triton fused), $\alpha$ corresponds to $m$. For our proposed method with memory-saving mode (Triton fused (mem save)), we can effectively regard $\alpha=2m$ because we store the real and imaginary parts of the state vector as bfloat16 for complex64 or float32 for complex128 to halve the memory required to store state vectors. Note that this memory-saving technique applies only to the output state vectors stored for each fused variational gate within an active checkpoint block. The input quantum states for checkpoint blocks are stored without type conversion because they are automatically managed by PyTorch’s checkpointing mechanism.

### 03 Apr 2026 v2

@@ Abstract
While real quantum devices have been increasingly used to conduct research focused on achieving quantum advantage or quantum utility in recent years, executing deep quantum circuits or performing quantum machine learning with large-scale data on current noisy intermediate-scale quantum devices remains challenging, making classical simulation essential for quantum machine learning research. However, such classical simulation often suffers from the cost of gradient calculations, requiring enormous memory or computational time. To address these problems, we propose a method to fuse multiple consecutive gates in each of the forward and backward paths to improve throughput by minimizing global memory accesses. As a result, we achieved approximately $20$ times throughput improvement for a Hardware-Efficient Ansatz with $12$ or more qubits, reaching over $30$ times improvement on a mid-range consumer GPU with limited memory bandwidth. By combining our proposed method with gradient checkpointing, we drastically reduced memory usage, making it possible to train a large-scale quantum machine learning model, a $20$-qubit, 1,0001{,}000-layer model with 60,00060{,}000 parameters, using 1,0001{,}000 samples in approximately $20$ minutes per epoch. This implies that we can train the model on large datasets, comprising tens of thousands of samples, like MNIST or CIFAR-10, within a realistic time frame (e.g., $20$ hours per epoch). Thus, our proposed method significantly accelerates such classical simulations, making a significant contribution to advancing research in quantum machine learning and variational quantum algorithms, such as verifying algorithms on large datasets or investigating learning theories of deep quantum circuits like barren plateaus.
@@ 1 Introduction
A popular approach for gradient calculation is the parameter-shift rule [mitarai2018quantum]. This approach is memory-efficient, but the number of entire quantum circuit executions scales linearly with the number of variational gates, making it impractical for large-scale models in classical simulation. On the other hand, the adjoint method is fast because the effective number of entire quantum circuit executions is independent of the number of variational gates. However, a practical implementation of the adjoint method faces a critical trade-off between memory usage and execution time regarding the management of intermediate states. A naive implementation stores all intermediate states in the forward path in order to reuse them in the backward path, causing memory usage to increase linearly with the number of gates. Alternatively, a memory-efficient implementation avoids storing intermediate states by sequentially applying each gate to recompute them during the backward path [jones2020efficient, luo2020yao], leading to a significant increase in computational time due to the repeated accesses to exponentially large quantum states [jones2020efficient, luo2020yao]. Therefore, existing simulators typically optimize only forward paths using gate fusion [suzuki2021qulacs, haner20175, smelyanskiy2016qhipster], or sequentially apply quantum gates to obtain intermediate states in the backward paths [jones2020efficient, luo2020yao], leaving the backward path bottlenecked by inefficient memory accesses.
@@
In this paper, we propose a method to fuse multiple consecutive quantum gates into a single fused operator for each of the forward and backward paths, without writing any intermediate states of the fused gates to global memory. This approach allows us to reduce the memory requirements for storing quantum states and decrease the number of accesses to global memory, leading to an enhancement of the arithmetic intensity. However, if we simply read the input state and temporarily store all recomputed intermediate states, the number of variables to be held in registers can easily exceed the register capacity. This causes register spilling and increases the computation time. Therefore, to reduce this register pressure, we minimize the number of intermediate states held in registers, and load unitary matrices whenever necessary, overwriting the previous values to reuse registers. As a result, our proposed method achieves approximately $20$ times improvement in throughput compared to a PyTorch native implementation for a typical single-layer Hardware-Efficient Ansatz (HEA) [kandala2017hardware]. In particular, for a mid-range GPU with limited memory bandwidth, we achieved over $30$ times improvement in throughput, highlighting the effectiveness of our proposed method on GPUs with limited memory bandwidth.
@@
Moreover, since our implementation uses Triton [triton2019] and PyTorch [Ansel_PyTorch_2_Faster_2024] and is integrated with PyTorch’s automatic differentiation, it is easy to utilize the existing PyTorch ecosystem. For example, by combining our proposed method with gradient checkpointing, we can reduce memory usage to $O(\sqrt{d})$ for $d$ HEA layers. In our numerical experiment, we performed a task of minimizing a sum of batched expectation values for an observable at the final states of a $20$-qubit with 1,0001{,}000 HEA layers (60,00060{,}000 parameters). As a result, we succeeded in training the model on 1,0001{,}000 samples in approximately $20$ minutes per epoch. This result indicates that we can train a model consisting of 60,00060{,}000 parameters using tens of thousands of samples within $20$ hours per epoch, i.e., a realistic time frame. Further acceleration of training can be easily realized using multiple GPUs with PyTorch’s distributed training utilities, such as DistributedDataParallel.
@@ 2.2 Adjoint method
Note that from Eq. (3), we need both the input quantum state $|\psi_{j-1}\rangle$ and the adjoint state $|\lambda_{j}\rangle$ to calculate the gradient $\frac{\partial L}{\partial\theta_{j}}$. This indicates that we need to preserve the input quantum states $|\psi_{j-1}\rangle$ to efficiently compute the gradients. Namely, a naive implementation requires $O(M\cdot 2^{n})$ memory to store all the input quantum states $\{|\psi_{0}\rangle,\dots,|\psi_{M-1}\rangle\}$ in memory. As the number of variational parameters $M$ grows, the required memory can easily exceed the memory capacity, making classical simulation infeasible. Even within the memory capacity, the computation time significantly increases due to repeated accesses to these exponentially large vectors in global memory. We will tackle these problems with our proposed method in the following sections.
@@ 3.1 Applying a quantum gate
As we mentioned in Sec. 2.1, an $n$-qubit state vector $|\psi\rangle$ is represented by a complex vector of length $2^{n}$. We describe $|\psi\rangle$ using computational basis states $|x\rangle$ and complex probability amplitudes $a_{x}$, whose index $x$ $(0\leq x<2^{n})$ corresponds to its binary representation $x_{n-1}\ldots x_{0}$ $(x=\sum_{i=0}^{n-1}x_{i}2^{i})$ of length $n$, as follows:
$|\psi\rangle=\sum_{x=0}^{2^{n}-1}a_{x}|x\rangle=a_{0\ldots 00}|0\cdots 00\rangle+a_{0\ldots 01}|0\cdots 01\rangle+\cdots+a_{1\ldots 10}|1\cdots 10\rangle+a_{1\ldots 11}|1\cdots 11\rangle$
@@
Next, we explain how to update amplitudes when applying two-qubit gates. First, when applying a CZ gate acting on control qubit $c$ and target qubit $t$ to a quantum state, we only need to multiply the amplitudes $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$) by 1-1, where the control and target bits of the index are both $1$. Similarly, when applying CNOT gate acting on control qubit $c$ and target qubit $t$ to a quantum state, we can update the quantum state by swapping the pair of amplitudes $a_{\ldots 1_{c}\ldots 0_{t}\ldots}$ and $a_{\ldots 1_{c}\ldots 1_{t}\ldots}$ (or $a_{\ldots 0_{t}\ldots 1_{c}\ldots}$ and $a_{\ldots 1_{t}\ldots 1_{c}\ldots}$), i.e., swapping the amplitudes corresponding to the $t$-th bit being 0 and $1$ for all indexes whrere the $c$-th bit is $1$.
@@ 3.2 Gate fusion in the forward path
Although the adjoint method requires the input state vector for each variational gate, our implementation stores only the output state vector of each fused variational gate during the forward path. As detailed in the next subsection, we then use this stored output state vector to recompute the input state vector for each constituent gate within the fused operator.
@@ 3.3 Gate fusion in the backward path
As shown in Eq. (3), when we calculate the gradients with respect to the variational parameters $\theta_{j+k}$, we need both the input quantum state $|\psi_{j+k-1}\rangle$ and the adjoint state $|\lambda_{j+k}\rangle$ for the corresponding variational gate. If we simply fuse $m$ consecutive variational gates in the backward path, it requires storing the elements of quantum states $|\psi_{j}\rangle,\ldots,|\psi_{j+m-1}\rangle$, $|\lambda_{j}\rangle,\ldots,|\lambda_{j+m}\rangle$ in registers. Therefore, as $m$ increases, the required memory can easily exceed the register capacity of the GPU, leading to register spilling into a low-speed memory region, which increases computation time.
To address this issue, by focusing on the reversibility of quantum gates, we recompute the intermediate states within GPU registers when needed, instead of storing all of them in global memory. Specifically, as shown in Fig. 1, given the quantum state $|\psi_{j+m}\rangle$ (stored during the forward path) and the adjoint state $|\lambda_{j+m}\rangle$ as inputs for the fused gates consisting of $m$ consecutive variational gates, each thread initially loads its assigned pairs of complex elements, $\left(a^{(j+m)}_{\dots 0_{t}\dots},a^{(j+m)}_{\dots 1_{t}\dots}\right)$ and $\left(r^{(j+m)}_{\dots 0_{t}\dots},r^{(j+m)}_{\dots 1_{t}\dots}\right)$, from global memory into its local registers, where $a^{(j+m)}_{x}$ represents the probability amplitude of the $(j+m)$-th quantum state $|\psi_{j+m}\rangle$ at index $x$ and $r^{(j+m)}_{x}$ represents the complex element of the $(j+m)$-th adjoint state $|\lambda_{j+m}\rangle$ at index $x$. Then, within a single fused GPU kernel, each thread performs the following procedure without accessing global memory to obtain intermediate states.
To recompute the input quantum states for $U_{j+k}(\theta_{j+k})$ and calculate the gradients $\frac{\partial L}{\partial\theta_{j+k}}$ for $k=1,\ldots,m$, we iterate the gate index of the fused gates in reverse order from $k=m$ to $k=1$. For each step $k$:
@@
    Each thread loads the elements of the unitary matrix $u_{j+k}(\theta_{j+k})$ and applies the matrix to its assigned pair of amplitudes of $|\psi_{j+k}\rangle$ in order to calculate the corresponding amplitudes of $|\psi_{j+k-1}\rangle=U_{j+k}^{\dagger}(\theta_{j+k})|\psi_{j+k}\rangle$:
    $\begin{pmatrix}a^{(j+k-1)}_{...0_{t}...}\\ a^{(j+k-1)}_{...1_{t}...}\end{pmatrix}=u_{j+k}^{\dagger}(\theta_{j+k})\begin{pmatrix}a^{(j+k)}_{...0_{t}...}\\ a^{(j+k)}_{...1_{t}...}\end{pmatrix}$,
    where $a^{(j+k)}_{x}$ represents the probability amplitude of the $(j+k)$-th intermediate state $|\psi_{j+k}\rangle$ at index $x$.
@@
    According to Eq. (3), each thread calculates its local contribution to the gradient $\frac{\partial L}{\partial\theta_{j+k}}$:
    $\text{Re}\left[\begin{pmatrix}r^{(j+k)*}_{...0_{t}...}&r^{(j+k)*}_{...1_{t}...}\end{pmatrix}\frac{\partial u_{j+k}(\theta_{j+k})}{\partial\theta_{j+k}}\begin{pmatrix}a^{(j+k-1)}_{...0_{t}...}\\ a^{(j+k-1)}_{...1_{t}...}\end{pmatrix}\right$],
    where $r^{(j+k)}_{x}$ represents the complex element of the $(j+k)$-th intermediate adjoint state $|\lambda_{j+k}\rangle$ at index $x$ and denotes the complex conjugate.
@@
    According to $|\lambda_{j+k-1}\rangle=U_{j+k}^{\dagger}(\theta_{j+k})|\lambda_{j+k}\rangle$, each thread updates its assigned pair of components of the adjoint state $|\lambda_{j+k}\rangle$ using the adjoint of the loaded gate $u_{j+k}(\theta_{j+k})$:
    $\begin{pmatrix}r^{(j+k-1)}_{...0_{t}...}\\ r^{(j+k-1)}_{...1_{t}...}\end{pmatrix}=u_{j+k}^{\dagger}(\theta_{j+k})\begin{pmatrix}r^{(j+k)}_{...0_{t}...}\\ r^{(j+k)}_{...1_{t}...}\end{pmatrix}$.
@@
Using this procedure, each thread only needs to hold a pair of probability amplitudes of the intermediate state $|\psi_{j+k-1}\rangle$ and a pair of complex elements of the adjoint state $|\lambda_{j+k}\rangle$, along with the elements of the unitary matrix $u_{j+k}(\theta_{j+k})$ and its derivative, in registers during kernel processing. To avoid register spilling, we load the elements of the unitary matrix $u_{j+k}(\theta_{j+k})$ and its derivative only when required, and reuse the registers by overwriting the previous elements of the state vector, the adjoint state, the unitary matrix, and its derivative. Compared to the method of storing all the input quantum states and adjoint states, although our method requires additional calculations to recompute the input quantum states for each $U_{j+k}(\theta_{j+k})$, it allows us to mitigate the bottleneck, i.e., minimize the accesses to global memory. This is expected to reduce the required memory and improve throughput, as fewer quantum states need to be stored.
@@ 3.4 Theoretical memory analysis with gradient checkpointing
To theoretically analyze the behavior of memory usage, we consider a typical deep parameterized quantum circuit consisting of a repeating layer structure, such as a Hardware-Efficient Ansatz (HEA). Let $d$ be the total number of HEA layers in a quantum circuit, and let $l_{\text{var}}$ and $l_{\text{const}}$ represent the number of variational gates and constant gates per HEA layer, respectively.
In the PyTorch native implementation of the adjoint method without gradient checkpointing, since we need to store the input quantum states of all gates for the backward path, the number of required state vectors to be stored is $(l_{\text{var}}+l_{\text{const}})d$. To mitigate this linear memory growth, we apply gradient checkpointing by dividing the quantum circuit into checkpoint blocks, each consisting of $b$ consecutive HEA layers. For simplicity, we assume that $d$ is divisible by $b$.
When combining the PyTorch native implementation with gradient checkpointing, the number of required state vectors to be stored for the backward path becomes $(l_{\text{var}}+l_{\text{const}})b+d/b$. The first term represents the stored intermediate states within a single active checkpoint block, and the second term ($d/b$) arises from the stored input quantum states for each checkpoint block, which are required to recalculate the intermediate state vectors within the checkpoint block during gradient checkpointing. Therefore, letting $M_{\text{sv}}$ be the required memory for a single state vector, the required memory $M_{total}$ (bytes) for the PyTorch native implementation is estimated as
@@
Similarly, as illustrated in Fig. 2, for our proposed method combined with gradient checkpointing, we must store $d/b$ input quantum states for the checkpoint blocks. Additionally, within the active checkpoint block, we only need to store $\left\lceil\frac{l_{\text{var}}}{m}\right\rceil\cdot b$ output state vectors of the fused variational gates, since the implementation of our proposed method does not require storing any input and output state vectors for constant gates. By introducing $\alpha$ as an effective number of fused gates to uniformly treat the memory reduction effects regardless of the type of saved tensors, the required memory (bytes) can be estimated as
@@
where the equality holds when $b^{*}=\sqrt{d/\left\lceil\frac{l_{\text{var}}}{\alpha}\right\rceil}$. Specifically, for our proposed method (Triton fused), $\alpha$ corresponds to $m$. For our proposed method with memory-saving mode (Triton fused (mem save)), we can effectively regard $\alpha=2m$ because we store the real and imaginary parts of the state vector as bfloat16 for complex64 or float32 for complex128 to halve the memory required to store state vectors. Note that this memory-saving technique applies only to the stored output state vectors of each fused variational gate within an active checkpoint block. The input quantum states for checkpoint blocks are stored without type conversion because they are automatically managed by PyTorch’s checkpointing mechanism.